OpenStudy (anonymous):

A rectangular piece of fabric is 16ft by 11ft. A triangular area with a height of 5.4ft and a base of 8.2ft is cut from the fabric. How much area is left over? A. 131.72 ft2 B. 153.86 ft2 C. 221.1 ft2 D. 263.44 ft2

OpenStudy (shane_b):

All you need to do is calculate the area of the rectangular and subtract off the area of the triangle.\[\Large A_{rect}=L*W\]\[\Large A_{triangle}=\frac{1}{2}b*h\]